OpenStudy (anonymous):

Solve the compound inequality 60 < 20u < 120. 3 < u < 6 6 > u > 7 6 < u < 3 3 > u < 6

OpenStudy (tkhunny):

Have you considered dividing everything by 20? Where did I get that number?

OpenStudy (anonymous):

the third is the answer right

OpenStudy (jdoe0001):

solve the inequality, thus you find out

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

the one thing to keep in mind is that, you have 3 terms, and any addition or multiplications will involve, not just a left-hand an rigth-hand side, but also a 3rd term

OpenStudy (anonymous):

i go 3 < u < 6

OpenStudy (anonymous):

yea thats what i got

OpenStudy (anonymous):

so the answer is first one

OpenStudy (anonymous):

okay can you please keep helping me...

OpenStudy (anonymous):

Solve the compound inequality 6b < 24 or 4b + 12 > 4. b < 4 or b > −2 b < 4 or b > 4 b < 4 or b > 2 b > 4 or b < −2

OpenStudy (anonymous):

what the answer

OpenStudy (anonymous):

Which of the following is an equivalent form of the compound inequality −44 > −2x − 8 >= −8?

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

−2x − 8 > −44 and −2x − 8 >= −8 −2x > −44 and −8 >= −8 −2x − 8 < −44 and −2x − 8 <= −8 −2x − 8 < −44 and −2x − 8 >= −8

OpenStudy (anonymous):

i hope you give me right answer please.... are you

OpenStudy (anonymous):

yea hold up let me check again

OpenStudy (anonymous):

yea the last one is the same way you did it